OpenStudy (anonymous):

Consider the equation 7x + 3y = 42. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points) Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points) @AriPotta ;* Sorry for bein' such a pain in da bum, but I need some help wife.

OpenStudy (anonymous):

solve for y first, isolate it

OpenStudy (anonymous):

for part 1 , you want it to look like y=mx +b

OpenStudy (anonymous):

Ok. One sec.

OpenStudy (anonymous):

So it would be: y = 2x + 14, right?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

not exactly. 3y=-7x + 42 What is the next step

OpenStudy (anonymous):

(This is why I need Harry/Ari) Oh holy christ, I didn't see the negative sign sorry.

OpenStudy (anonymous):

It would be y= -2x + 14 right? Have a feeling that's still wrong o.o Math & I don't get along.

OpenStudy (aripotta):

sorry, i didn't notice the notification! </3

OpenStudy (anonymous):

think of it like this if I had 5+2 = 7 and I told you to find the 2 by itself you would move the 5 over by subtracting, right? so 2=7-5

OpenStudy (anonymous):

its the same thing except you have variables

OpenStudy (anonymous):

Harry; you left me. Zebo; mmk I think I see what you're saying.

OpenStudy (aripotta):

i'm sorry! :c

OpenStudy (anonymous):

We're over, clearly. Unless you can explain this too me? :*

OpenStudy (aripotta):

ok so you've gotten this far: 3y = -7x + 42 but it's still not in slope-intercept form. you gotta divide both sides by 3. what do you get when you do that?

OpenStudy (anonymous):

Eh, sorry. Computer is dreadfully slow tonight. I got ; y = -2x + 14 but that isn't right, is it Harry?

OpenStudy (aripotta):

no because -7/3 isn't -2

OpenStudy (aripotta):

you're going to keep it -7/3 y = -7/3x + 14

OpenStudy (anonymous):

Oh I see what you did there, Harry.

OpenStudy (aripotta):

so when graphing, you plot the y intercept first. which is (0,14)

OpenStudy (aripotta):

then you use the slope to plot the next point. slope is rise over run, so from the y-intercept, go down 7 and over to the right 3

OpenStudy (anonymous):

Ya, I see what you're gettin' at Harry :3

OpenStudy (aripotta):

plot that, then put a line through le points

OpenStudy (aripotta):

and you've just graphed teh line

OpenStudy (anonymous):

*le'points. Thanks Harry :3 Once again you've saved Hogwarts I mean my booty.

OpenStudy (aripotta):

well that's because i love ur booty

OpenStudy (aripotta):

oh we have to do the interepts method too

OpenStudy (aripotta):

intercepts*

OpenStudy (anonymous):

Well yes of course you love my booty, and Cols. *duuuh* Oh yeah.

OpenStudy (aripotta):

so this time, we gotta find the x- and y-intercepts. we'll use the standard form. 7x + 3y = 42 for x-intercept, substitute 0 for y 7x + 3(0) = 42 7x = 42 x = 6 x-intercept = (6,0) for the y-intercept, substitute 0 for x 7(0) + 3y = 42 3y = 42 y = 14 y-intercept = (0,14)

OpenStudy (aripotta):

now just plot those 2 points and put a line through em

OpenStudy (anonymous):

Oh I see what you did there, Harry. You're to smart for your own good, Potter. This exam has taken me like 2 hours, because Tumblr started working again right as I sat down to do it. >.<

OpenStudy (aripotta):

lol

OpenStudy (aripotta):

how many questions left? or are you done?

OpenStudy (anonymous):

After this one, like ... 2 I think. So distracted by Eddie in a suit, Ari... I can't take it anymore. Anyways, almost done - then we can draw faces on cats together. Yay!

OpenStudy (aripotta):

yey! we still have the twiddla open so pop in once you're done, mmk

OpenStudy (anonymous):

mmk, Harry I will.